Kyoto Joe, Inc., sells earnings forecasts for Japanese securities. Its credit terms are 1/20, net 40. Based on experience, 65 percent of all customers will take the discount.

a.

What is the average collection period for the company? (Use 365 days a year. Do not round intermediate calculations and round your answer to 2 deceimal places, e.g., 32.16.)

b. If the company sells 1,450 forecasts every month at a price of $1,240 each, what is its average balance sheet amount in accounts receivable?
Answer #1

a. Average collection period = 0.65(20 days) + 0.35(40 days) = 27 days

b. Average balance = 1,450($1,240)(27)(12/365) = $1,596,032.88